OpenStudy (anonymous):

What is the derivative of y=[(1+sinx)/(1+cosx)]^2

OpenStudy (anonymous):

\[[\frac{f}{g}]'=\frac{ f'g -fg'}{ g^2 }\]

OpenStudy (anonymous):

first use chain rule to drop the 2

OpenStudy (anonymous):

Use the product rule for this. Well that's what I would use.

OpenStudy (anonymous):

Actually, both methods are good.

OpenStudy (rob1525):

OMG! Do you want the answer because its really long.

OpenStudy (anonymous):

I think I got it, I used the chain rule, and then the quotient rule for the g'(x) portion of the chain rule. This is a very, very ugly problem.

OpenStudy (anonymous):

This is what I got. You can use and compare you your answer. 

OpenStudy (anonymous):

\[2(\frac{\cos x(1+\cos x)+\sin x(1+\sin x)}{(\cos x +1)^2})\]

OpenStudy (anonymous):

not that bad

OpenStudy (rob1525):

wow theres a lot of ways to do this problem. I got \[(2*(\sin(x) +1)*(\cos(x) +\sin(x) +1)/(\cos(x)+1)^3\]

OpenStudy (rob1525):

or (2(5cos(x)+5sin(x)cos(x)+3cos^(2)(x)+3sin(x)cos^(2)(x)+sin(x)+2sin^(2)(x)+sin^(3)(x)))/(1+cos(x))